OpenStudy (anonymous):

MEDAL & FAN ANYBODY ? A car drives 8 miles south on a freeway, then exits and travels west for 6 miles. How far is the car from its starting position? a. 9 miles b. 10 miles c. 11 miles d. 12 miles

OpenStudy (studygurl14):

Use the Pythagorean theorem. Draw a triangle to help you

OpenStudy (anonymous):

I think it's D

OpenStudy (studygurl14):



OpenStudy (anonymous):

@StudyGurl14 would it be 10 ?

OpenStudy (studygurl14):

How did you get D. (PS. it is the incorrect answer)

OpenStudy (studygurl14):

Pythagorean Theorem: \[a^2+b^2=c^2\]

OpenStudy (chosenmatt):

one sec

OpenStudy (studygurl14):

Yes! Good job!.

OpenStudy (studygurl14):

10 is the correct answer, because... \[8^2+6^2=c^2\] \[64+36=c^2\] \[100=c^2 \rightarrow c = \sqrt{100} = 10\]

OpenStudy (anonymous):

have you heard of the three four five triangle the answer is simple divide each by 2 you get two points of the triangle and then you know that the answer is 10